OpenStudy (anonymous):

Write an equation of the line containing the given point and parallel to the given line. (2, -7); 5x-8y=7

OpenStudy (campbell_st):

ok, what is the slope of the given line..?

OpenStudy (anonymous):

The slope is -5x/8

OpenStudy (campbell_st):

close 5x - 7 = 8y so y = 5/8 x - 7 so m = 5/8 and the point is (2, -7) an easy form of the line is y = mx + b or \[y = \frac{5x}{8} + b\] to find b, substitute the point, x = 2 and y = -7 hope it helps...

OpenStudy (campbell_st):

they you can re-write it in standard form...

OpenStudy (anonymous):

Thanks Can you help me to write it in standard form?

OpenStudy (campbell_st):

ok... so what did you get for b?

OpenStudy (anonymous):

-7

OpenStudy (campbell_st):

not quite, if you substitute you get \[-7 = \frac{5 \times 2}{8} + b\] or \[-\frac{66}{8} = b\] so in slope intercept form the equation is \[y = \frac{5}{8} x - \frac{66}{8}\] now multiply every term by 8 \[8y = 5x - 66\] so rewriting it in standard form 5x - 8y = 66

OpenStudy (campbell_st):

lol... oops... just realised the easy method if you let 5x - 8y = b then substitute x = 2 and y = -7 you get 5x - 8y = 66 sorry, just a bit slow this morning

OpenStudy (anonymous):

Okay Thank you so much